OpenStudy (anonymous):

Find the equation of the tangent line to the gven curve at the given point y=2sinx at (pi/6, 1)

OpenStudy (anonymous):

Am i suppose to use unit circle..?

OpenStudy (anonymous):

Tangent line at ( pi/6, 1) is: y - 1 = m (x - pi/6 ) m = f'( pi /6) Can you do it?

OpenStudy (anonymous):

do i u use the unit circle for pi/6? i tried it without using unit circle, and get some really weird answer

OpenStudy (p0sitr0n):

f`(number) = 0 ?

OpenStudy (anonymous):

f' for this equation would be, if I'm not mistaken 2cos(x)

OpenStudy (anonymous):

yea i think f' is 2cos (x) too

OpenStudy (anonymous):

as for your question if you use the unit circle. yea; I do believe you use the value of cos at pi/6

OpenStudy (anonymous):

then i will get critical numbers using unit circle what do i do with it?

OpenStudy (anonymous):

f' (x) = 2 cosx -> f' ( pi/6 ) = .....

OpenStudy (anonymous):

Chlorophyll is right, use the equation he gave above and plug in what you know already. That should give you the answer. y-1=m(x-pi/6) where m=2cos(x)